Sarah Meeham blends coffee for Tasti-Delight. She needs to prepare 190 pounds of blended coffee beans selling for $4.55 per pound. She plans to do this by blending together a high-quality bean costing $5.50 per pound and a cheaper bean at $3.25 per pound. To the nearest pound, find how much high-quality coffee bean and how much cheaper coffee bean she should blend.
Answer by rwm(914)

h+c=190
5.50h+3.25c=190*4.55
c=about 80
h=about 110